# SU(3) gauge field of magnons in antiferromagnetic skyrmion crystals

### v2

@@ Abstract
for $\ell\neq\ell^{\prime}$ with unit and Pauli matrices $\tau^{\mu}$ ($\mu=0,x,y,z$) and Levi-Civita symbol ϵ′′subscriptitalic-$\epsilon_{\ell\ell^{\prime}\ell^{\prime\prime}}$. Here, we introduce four vector fields as $A_{\ell,\mu}(\bm{r})=i\bm{e}_{\ell}^{+}(\bm{r})\cdot\partial_{\mu}\bm{e}_{\ell}^{-}(\bm{r})$ and $\xi_{\mu}(\bm{r})=(1/3)[\bm{m}_{1}(\bm{r})\cdot\partial_{\mu}\bm{m}_{2}(\bm{r})+\bm{m}_{2}(\bm{r})\cdot\partial_{\mu}\bm{m}_{3}(\bm{r})+\bm{m}_{3}(\bm{r})\cdot\partial_{\mu}\bm{m}_{1}(\bm{r})]$ ($\mu=x,y$). In particular, $\bm{A}_{\ell}(\bm{r})$ can be interpreted as the U(1) gauge field arises from the FM-SkX on the sublattice $\ell$.
The effective magnon Hamiltonian (7) is complicated and conceals its gauge structure. To elucidate the hidden gauge structure, we introduce new bosonic operators $\hat{\gamma}_{n}(\bm{r})$ ($n=1,2,3$) as the linear combination of $\hat{b}_{\ell}(\bm{r})$ and $\hat{b}_{\ell}^{\dagger}(\bm{r})$ such that the resulting effective Hamiltonian has the following two properties; (i) the conventional $120^{\circ}$ order, namely spatially uniform $\bm{m}_{\ell}(\bm{r})$, gives rise to three decoupled magnons with well-known linear dispersions and (ii) the equation of motion for $\hat{\gamma}_{n}(\bm{r})$ is relativistic, which is expected in antiferromagnets. We find the Bogoliubov transformation that satisfies these properties as
@@
^$\hat{\mathcal{H}}_{\mathrm{eff}}^{\mathrm{AFM}}$ $\simeq\frac{1}{2}\int d^{2}\bm{r}\ \hat{\Psi}^{\dagger}(\bm{r})H(\bm{r})\hat{\Psi}(\bm{r}),$ (12)
@@
$H(\bm{r})$ $=\frac{3\sqrt{2}}{8}JSa^{2}\left[\tilde{I}\otimes\tau^{x}\nabla^{2}-i\bm{T}(\bm{r})\otimes(\tau^{0}+3\tau^{x})\cdot\bm{\nabla}\right]$
$\hskip 20.0pt+\frac{9\sqrt{2}}{4}JSI\otimes(\tau^{0}+\tau^{x}),$ (13)
with $\tilde{I}=\mathrm{diag}(2,1,1)$ and $I=\mathrm{diag}(1,1,1)$. Here, we ignore terms that give higher-order contributions to the equation of motion for $\hat{\gamma}_{n}(\bm{r})$. The difference $\tilde{I}\neq I$ reflects the fact that one spin-wave mode has a different velocity in the $120^{\circ}$ order phase of the triangular-lattice antiferromagnet [47]. For simplicity, we neglect this difference and replace $\tilde{I}$ with $I$ in the following. The $3\times 3$ matrix $T_{\mu}(\bm{r})$ ($\mu=x,y$) is given by $\bm{T}(\bm{r})=\bm{T}^{(2)}(\bm{r})\lambda_{2}+\bm{T}^{(5)}(\bm{r})\lambda_{5}$, where
$\bm{T}^{(2)}(\bm{r})$ $=\frac{1}{3}\sum_{\ell}\bm{A}_{\ell}(\bm{r})\cos\theta_{\ell},$ (14)
$\bm{T}^{(5)}(\bm{r})$ $=\frac{1}{3}\sum_{\ell}\bm{A}_{\ell}(\bm{r})\sin\theta_{\ell},$ (15)
and $\lambda_{\alpha}$ ($\alpha=1,2,\ldots,8$) are Gell-Mann matrices
$\lambda_{2}=\begin{pmatrix}0&-i&0\\ i&0&0\\ 0&0&0\end{pmatrix},\hskip 10.0pt\lambda_{5}=\begin{pmatrix}0&0&-i\\ 0&0&0\\ i&0&0\end{pmatrix},\hskip 10.0pt\lambda_{7}=\begin{pmatrix}0&0&0\\ 0&0&-i\\ 0&i&0\end{pmatrix}.$ (16)
Since $T_{\mu}(\bm{r})$ is Hermitian and traceless, it belongs to the Lie algebra of the SU(3) group.
@@
The role of $\bm{T}(\bm{r})$ becomes more apparent upon deriving the equation of motion for $\hat{\bm{\gamma}}(\bm{r})=(\hat{\gamma}_{1}(\bm{r}),\hat{\gamma}_{2}(\bm{r}),\hat{\gamma}_{3}(\bm{r}))^{T}$. From the effective Hamiltonian (12), we obtain the equation of motion for $\hat{\Psi}(\bm{r},t)$ as $i\hbar\partial_{t}\hat{\Psi}(\bm{r},t)=(I\otimes\tau^{z})H(\bm{r})\hat{\Psi}(\bm{r},t)$. Then the equation of motion for $\hat{\bm{\gamma}}(\bm{r},t)$ is derived as
@@
$\hbar^{2}\frac{\partial^{2}}{\partial t^{2}}\hat{\bm{\gamma}}(\bm{r},t)=\frac{27}{8}(JSa)^{2}\left(\bm{\nabla}I-i\bm{T}(\bm{r})\right)^{2}\hat{\bm{\gamma}}(\bm{r},t).$ (17)
For the conventional $120^{\circ}$ order, we have $\bm{T}(\bm{r})=\bm{0}$ and the low-energy excitations are described by three independent magnons with linear dispersion $\varepsilon(\bm{k})\simeq(3/2)^{3/2}JSa|\bm{k}|$. For the AFM-SkXs, the real-space spin texture induces finite $\bm{T}(\bm{r})$, which couples the three species of magnons with their kinetic motion and can be interpreted as the SU(3) gauge field.
We remark that when $T_{x}(\bm{r})$ and $T_{y}(\bm{r})$ commute, there exists the unitary matrix $U(\bm{r})$ that diagonalizes $T_{x}(\bm{r})$ and $T_{y}(\bm{r})$ simultaneously, and the equation of motion (17) can be decomposed into three independent magnons with U(1) gauge fields as
$\hbar^{2}\frac{\partial^{2}}{\partial t^{2}}\hat{\gamma}_{n}^{\prime}(\bm{r},t)=\frac{27}{8}(JSa)^{2}\left(\bm{\nabla}-i\bm{A}_{n}^{\prime}(\bm{r})\right)^{2}\hat{\gamma}_{n}^{\prime}(\bm{r},t),$ (18)
where $\hat{\gamma}_{n}^{\prime}(\bm{r})=\sum_{m=1}^{3}U_{m,n}^{*}(\bm{r})\hat{\gamma}_{m}(\bm{r})$ ($n=1,2,3$) and $\bm{A}_{n}^{\prime}(\bm{r})$ is the U(1) gauge field determined by the eigenvalues of $T_{\mu}(\bm{r})$ ($\mu=x,y$). Since $T_{\mu}(\bm{r})$ is the traceless matrix, the U(1) gauge fields satisfy $\sum_{n}\bm{A}_{n}^{\prime}(\bm{r})=\bm{0}$ and the associated magnetic fields $B_{n}^{\prime}(\bm{r})=\partial_{x}A_{n,y}^{\prime}(\bm{r})-\partial_{y}A_{n,x}^{\prime}(\bm{r})$ cancel out each other, $\sum_{n}B_{n}^{\prime}(\bm{r})=0$. This cancellation indicates that the net thermal Hall conductivity from three species reduces to zero. Therefore, $\bm{T}(\bm{r})$ with $[T_{x}(\bm{r}),T_{y}(\bm{r})]=0$ does not contribute to the magnon thermal Hall effect.
@@
Next, we consider the effect of finite $[T_{x}(\bm{r}),T_{y}(\bm{r})]$. To this end, we focus on the field strength defined as $F_{xy}(\bm{r})=\partial_{x}T_{y}(\bm{r})-\partial_{y}T_{x}(\bm{r})-i[T_{x}(\bm{r}),T_{y}(\bm{r})]$, which is the counterpart of the fictitious magnetic field in the U(1) gauge field. Since $F_{xy}(\bm{r})$ belongs to the Lie algebra of the SU(3) group, it can be expanded as $F_{xy}(\bm{r})=\sum_{\alpha=1}^{8}F_{xy}^{(\alpha)}(\bm{r})\lambda_{\alpha}$. The field strength is characterized by the eight real values $F_{xy}^{(\alpha)}(\bm{r})$, whose nonzero elements are calculated from Eqs. (14) and (15) as
$F_{xy}^{(2)}(\bm{r})$ $=\frac{1}{3}\sum_{\ell}\cos\theta_{\ell}B_{\ell}(\bm{r}),$ (19)
$F_{xy}^{(5)}(\bm{r})$ $=\frac{1}{3}\sum_{\ell}\sin\theta_{\ell}B_{\ell}(\bm{r}),$ (20)
$F_{xy}^{(7)}(\bm{r})$ $=\frac{1}{4}\bm{n}(\bm{r})\cdot(\partial_{x}\bm{n}(\bm{r})\times\partial_{y}\bm{n}(\bm{r})),$ (21)
@@
where $B_{\ell}(\bm{r})=\bm{m}_{\ell}(\bm{r})\cdot(\partial_{x}\bm{m}_{\ell}(\bm{r})\times\partial_{y}\bm{m}_{\ell}(\bm{r}))$ is the fictitious magnetic field on the sublattice $\ell$. In particular, $F_{xy}^{(7)}(\bm{r})$ comes from the commutation relation $[T_{x}(\bm{r}),T_{y}(\bm{r})]$ and is determined by the scalar chirality of the vector field $\bm{n}(\bm{r})$.
@@
In analogy to the case of the FM-SkXs, we focus on the uniform elements of the field strength by replacing $F_{xy}^{(\alpha)}(\bm{r})$ with $\bar{F}_{xy}^{(\alpha)}=(1/V)\int d^{2}\bm{r}\ F_{xy}^{(\alpha)}(\bm{r})$. $\bar{F}_{xy}^{(2)}$ and $\bar{F}_{xy}^{(5)}$ are written in terms of the average fictitious magnetic field $\bar{B}_{\ell}=(1/V)\int d^{2}\bm{r}\ B_{\ell}(\bm{r})$. We also assume $\bar{B}_{1}=\bar{B}_{2}=\bar{B}_{3}$ since three FM-SkXs in the AFM-SkXs typically have the same skyrmion density [39, 40, 41, 42]. Within this approximation, $\bar{F}_{xy}^{(2)}$ and $\bar{F}_{xy}^{(5)}$ are reduced to zero, whereas $\bar{F}_{xy}^{(7)}$ takes the finite value which is proportional to the skyrmion density of the vector field $\bm{n}(\bm{r})$ defined as $(1/4\pi V)\int d^{2}\bm{r}\ \bm{n}(\bm{r})\cdot(\partial_{x}\bm{n}(\bm{r})\times\partial_{y}\bm{n}(\bm{r}))$. Returning to the original lattice, finite $\bar{F}_{xy}^{(7)}$ generates a uniformly distributed flux $\Phi=\bar{F}_{xy}^{(7)}\lambda_{7}v^{\prime}/2$ as shown in Fig. 1(d), which breaks the effective time-reversal symmetry and contributes to the magnon thermal Hall effect.

### v3

@@ Abstract
for $\ell\neq\ell^{\prime}$ with unit and Pauli matrices $\tau^{\nu}$ ($\nu=0,x,y,z$) and Levi-Civita symbol $\epsilon_{\ell\ell^{\prime}\ell^{\prime\prime}}$. Here, we introduce four vector fields as $A_{\ell,x}(\bm{r})=i\bm{e}_{\ell}^{+}(\bm{r})\cdot\partial_{x}\bm{e}_{\ell}^{-}(\bm{r})+(2D/Ja)m_{\ell}^{y}(\bm{r})$, $A_{\ell,y}(\bm{r})=i\bm{e}_{\ell}^{+}(\bm{r})\cdot\partial_{y}\bm{e}_{\ell}^{-}(\bm{r})-(2D/Ja)m_{\ell}^{x}(\bm{r})$, and $\xi_{\mu}(\bm{r})=(1/3)[\bm{m}_{1}(\bm{r})\cdot\partial_{\mu}\bm{m}_{2}(\bm{r})+\bm{m}_{2}(\bm{r})\cdot\partial_{\mu}\bm{m}_{3}(\bm{r})+\bm{m}_{3}(\bm{r})\cdot\partial_{\mu}\bm{m}_{1}(\bm{r})]$ ($\mu=x,y$). In particular, $\bm{A}_{\ell}(\bm{r})$ can be interpreted as the U(1) gauge field on the sublattice $\ell$.
Figure 2: Skyrmion crystals and gauge fields of magnons. (a), (b) Schematic illustration of the real-space spin configuration of the (a) FM-SkXs and (b) AFM-SkXs on the triangular lattice. The AFM-SkXs consist of three intertwined FM-SkXs as shown in different colors, giving rise to the three-sublattice structure. (c), (d) Schematic illustration of the uniform flux in the low-energy magnon systems for the (c) FM-SkXs and (d) AFM-SkXs. The clockwise (counterclockwise) arrows indicate the negative (positive) direction of the flux. The three different colors on each site in (d) represent the three internal degrees of freedom for the magnons. The uniform flux $\Phi$ originates from the commutation relation of an SU(3) gauge field. (e) Schematic illustration of the triangular-lattice ferromagnet with the DM interaction $\bm{D}_{i,j}\parallel\bm{e}^{Z}$. Gray arrows indicate the order of spins in the DM term. (f) Staggered flux pattern generated by the DM interaction. The system has the $\pi$-rotation symmetry along any straight line in the triangular lattice, which prohibits the finite thermal Hall conductivity.
The effective magnon Hamiltonian (8) is complicated and conceals its gauge structure. To elucidate the hidden gauge structure, we introduce new bosonic operators $\hat{\gamma}_{m}(\bm{r})$ ($m=1,2,3$) as the linear combination of $\hat{b}_{\ell}(\bm{r})$ and $\hat{b}_{\ell}^{\dagger}(\bm{r})$ such that the resulting effective Hamiltonian has the following two properties; (i) the conventional $120^{\circ}$ order, namely spatially uniform $\bm{m}_{\ell}(\bm{r})$, gives rise to three decoupled magnons with well-known linear dispersions and (ii) the equation of motion for $\hat{\gamma}_{m}(\bm{r})$ is relativistic, which is expected in antiferromagnets. We find the Bogoliubov transformation that satisfies these properties as
@@
$\hat{\mathcal{H}}_{\mathrm{eff}}^{\mathrm{AFM}}$ $\simeq\frac{1}{2}\int\mathrm{d}^{2}\bm{r}\ \hat{\Psi}^{\dagger}(\bm{r})H(\bm{r})\hat{\Psi}(\bm{r}),$ (13)
@@
$H(\bm{r})$ $=\frac{3\sqrt{2}}{8}JSa^{2}\left[\tilde{I}\otimes\tau^{x}\nabla^{2}-i\bm{G}(\bm{r})\cdot\bm{\nabla}\right]$
$\hskip 20.0pt+\frac{9\sqrt{2}}{4}JSI\otimes(\tau^{0}+\tau^{x})+g\mu_{\mathrm{B}}BM,$ (14)
with $\tilde{I}=\mathrm{diag}(2,1,1)$. The $6\times 6$ matrices $\bm{G}(\bm{r})=(G_{x}(\bm{r}),G_{y}(\bm{r}))$ and $M$ are defined as
$\bm{G}(\bm{r})$ $=\bm{A}_{\mathrm{c}}(\bm{r})\left[\lambda_{2}\otimes(\tau^{0}+3\tau^{x})+\frac{1}{\sqrt{2}}(\lambda_{3}-\sqrt{3}\lambda_{8})\otimes\tau^{z}\right]$
$\quad+\bm{A}_{\mathrm{s}}(\bm{r})\left[\lambda_{5}\otimes(\tau^{0}+3\tau^{x})+\sqrt{2}\lambda_{6}\otimes\tau^{z}\right]$
$\quad-\sqrt{3}\bm{\xi}(\bm{r})\lambda_{7}\otimes(\tau^{0}+\tau^{x}),$ (15)
$M$ $=\frac{\sqrt{2}}{12}m_{\mathrm{u}}^{z}\left[I\otimes(9\tau^{0}+\tau^{x})-\sqrt{3}(\sqrt{3}\lambda_{3}+\lambda_{8})\otimes\tau^{x}\right],$ (16)
where $\bm{A}_{\mathrm{c}}(\bm{r})=(1/3)\sum_{\ell}\bm{A}_{\ell}(\bm{r})\cos\theta_{\ell}$, $\bm{A}_{\mathrm{s}}(\bm{r})=(1/3)\sum_{\ell}\bm{A}_{\ell}(\bm{r})\sin\theta_{\ell}$, and $\lambda_{\alpha}$ ($\alpha=1,2,\ldots,8$) are the Gell-Mann matrices. Here, we ignore terms proportional to $\sum_{\ell}\bm{A}_{\ell}(\bm{r})\simeq 0$, $m_{\mathrm{c}}^{z}(\bm{r})=(1/3)\sum_{\ell}m_{\ell}^{z}(\bm{r})\cos\theta_{\ell}$, and $m_{\mathrm{s}}^{z}(\bm{r})=(1/3)\sum_{\ell}m_{\ell}^{z}(\bm{r})\sin\theta_{\ell}$. The latter two values vanish upon spatial averaging and only give higher-order contributions to the equation of motion for $\hat{\gamma}_{m}(\bm{r})$.
@@
The gauge structure becomes more apparent upon deriving the equation of motion for $\hat{\bm{\gamma}}(\bm{r})=(\hat{\gamma}_{1}(\bm{r}),\hat{\gamma}_{2}(\bm{r}),\hat{\gamma}_{3}(\bm{r}))^{T}$. From the effective Hamiltonian (13), we obtain the equation of motion for $\hat{\Psi}(\bm{r},t)$ as $i\hbar\partial_{t}\hat{\Psi}(\bm{r},t)=(I\otimes\tau^{z})H(\bm{r})\hat{\Psi}(\bm{r},t)$. Then the equation of motion for $\hat{\bm{\gamma}}(\bm{r},t)$ is derived as [51]
@@
$\hbar^{2}\frac{\partial^{2}}{\partial t^{2}}\hat{\bm{\gamma}}(\bm{r},t)=\left[\frac{27}{8}(JSa)^{2}\tilde{I}\left(\bm{\nabla}I-i\bm{T}(\bm{r})\right)^{2}+\Xi\right]\hat{\bm{\gamma}}(\bm{r},t),$ (17)
where the $3\times 3$ matrices $\bm{T}(\bm{r})=(T_{x}(\bm{r}),T_{y}(\bm{r}))$ and $\Xi$ are defined as
$\tilde{I}\bm{T}(\bm{r})$ $=\bm{A}_{\mathrm{c}}(\bm{r})\left[\lambda_{2}-\frac{\sqrt{2}}{4}(\lambda_{3}-\sqrt{3}\lambda_{8})\right]+\bm{A}_{\mathrm{s}}(\bm{r})\left(\lambda_{5}-\frac{\lambda_{6}}{\sqrt{2}}\right)$ (18)
$\Xi$ $=\frac{3}{4}JSg\mu_{\mathrm{B}}Bm_{\mathrm{u}}^{z}\left[8I+\sqrt{3}(\sqrt{3}\lambda_{3}+\lambda_{8})\right].$ (19)
Since $T_{\mu}(\bm{r})$ is Hermitian and traceless, it belongs to the Lie algebra of the SU(3) group and can be interpreted as the SU(3) gauge field. The diagonal matrix $\Xi$ describes the potential term from the coupling between the magnetic field and uniform magnetization. For the conventional $120^{\circ}$ order, namely for spatially-independent $\bm{m}_{\ell}(\bm{r})$ with $m_{\mathrm{u}}^{z}=0$, we have $\bm{T}(\bm{r})=\bm{0}$ and $\Xi=0$, reproducing three independent magnons with linear dispersions $\varepsilon_{1}(\bm{k})\simeq\sqrt{2}(3/2)^{3/2}JSa|\bm{k}|$ and $\varepsilon_{2}(\bm{k})=\varepsilon_{3}(\bm{k})\simeq(3/2)^{3/2}JSa|\bm{k}|$ [52].
@@
The SU(3) gauge field $\bm{T}(\bm{r})$ generates a physical field, called field strength, defined as $F_{xy}(\bm{r})=\partial_{x}T_{y}(\bm{r})-\partial_{y}T_{x}(\bm{r})-i[T_{x}(\bm{r}),T_{y}(\bm{r})]$, which is the counterpart of the fictitious magnetic field in the U(1) gauge field. Since $\tilde{I}F_{xy}(\bm{r})$ belongs to the Lie algebra of the SU(3) group, it can be expanded as $\tilde{I}F_{xy}(\bm{r})=\sum_{\alpha=1}^{8}\tilde{F}_{xy}^{(\alpha)}(\bm{r})\lambda_{\alpha}$. The field strength is characterized by the eight real values $\tilde{F}_{xy}^{(\alpha)}(\bm{r})$, whose nonzero elements are calculated from Eq. (18) as
$\tilde{F}_{xy}^{(2)}(\bm{r})=B_{\mathrm{c}}(\bm{r}),\hskip 5.0pt\tilde{F}_{xy}^{(3)}(\bm{r})=-\frac{\sqrt{2}}{4}B_{\mathrm{c}}(\bm{r}),\hskip 5.0pt\tilde{F}_{xy}^{(8)}(\bm{r})=\frac{\sqrt{6}}{4}B_{\mathrm{c}}(\bm{r})$ (20)
$\tilde{F}_{xy}^{(5)}(\bm{r})=B_{\mathrm{s}}(\bm{r}),\hskip 5.0pt\tilde{F}_{xy}^{(6)}(\bm{r})=-\frac{B_{\mathrm{s}}(\bm{r})}{\sqrt{2}}$ (21)
$\tilde{F}_{xy}^{(7)}(\bm{r})=-\frac{1}{8}\bm{n}(\bm{r})\cdot(\partial_{x}\bm{n}(\bm{r})\times\partial_{y}\bm{n}(\bm{r}))$
$\hskip 37.5pt-\frac{1}{3}\frac{D}{Ja}\left([\bm{n}(\bm{r})\times\partial_{x}\bm{n}(\bm{r})]^{x}+[\bm{n}(\bm{r})\times\partial_{y}\bm{n}(\bm{r})]^{y}\right)$
$\hskip 37.5pt-\frac{\sqrt{3}}{12}\left(\frac{2D}{Ja}\right)^{2}n^{z}(\bm{r}),$ (22)
@@
where $B_{\mathrm{c}}(\bm{r})=(1/3)\sum_{\ell}B_{\ell}(\bm{r})\cos\theta_{\ell}$, $B_{\mathrm{s}}(\bm{r})=(1/3)\sum_{\ell}B_{\ell}(\bm{r})\sin\theta_{\ell}$, and $B_{\ell}(\bm{r})=\bm{m}_{\ell}(\bm{r})\cdot(\partial_{x}\bm{m}_{\ell}(\bm{r})\times\partial_{y}\bm{m}_{\ell}(\bm{r}))-(2D/Ja)\sum_{\mu=x,y}\partial_{\mu}m_{\ell}^{\mu}(\bm{r})$ is the fictitious magnetic field on the sublattice $\ell$. In particular, $\tilde{F}_{xy}^{(7)}(\bm{r})$ comes from the commutation relation $T_{x}(\bm{r}),T_{y}(\bm{r})]$ and is determined by the scalar chirality of the vector field $\bm{n}(\bm{r})$ and DM interaction.
@@
In analogy to the case of the FM-SkXs, we focus on the uniform elements of the field strength by replacing $\tilde{F}_{xy}^{(\alpha)}(\bm{r})$ with $\bar{F}_{xy}^{(\alpha)}=(1/V)\int\mathrm{d}^{2}\bm{r}\ \tilde{F}_{xy}^{(\alpha)}(\bm{r})$. We assume $\bar{B}_{1}=\bar{B}_{2}=\bar{B}_{3}$ since three FM-SkXs in the AFM-SkXs typically have the same skyrmion density [39, 40, 41, 42]. Within this approximation, the terms proportional to $B_{\mathrm{c}}(\bm{r})$ and $B_{\mathrm{s}}(\bm{r})$ are reduced to zero, whereas $\bar{F}_{xy}^{(7)}$ takes the finite value. The first term in $\bar{F}_{xy}^{(7)}$ is proportional to the skyrmion density of the vector field $\bm{n}(\bm{r})$ defined as $\rho_{\mathrm{AFM-SkX}}=(1/4\pi V)\int\mathrm{d}^{2}\bm{r}\ \bm{n}(\bm{r})\cdot(\partial_{x}\bm{n}(\bm{r})\times\partial_{y}\bm{n}(\bm{r}))$, and the other two terms are from the coupling between the DM interaction and vector field $\bm{n}(\bm{r})$. Returning to the original lattice, finite $F_{xy}(\bm{r})$ generates a uniformly distributed flux $\Phi=\bar{F}_{xy}^{(7)}\lambda_{7}v^{\prime}/2$ as shown in Fig. 2(d), which breaks the effective time-reversal symmetry and contributes to the magnon thermal Hall effect.